OpenStudy (anonymous):

2 theta is in what quadrant?

OpenStudy (whpalmer4):

That depends on the value of \(\theta\), of course...

OpenStudy (anonymous):

the problem does not tell me the value. What i asked is what I was asked :o

OpenStudy (whpalmer4):

There must be other context. Otherwise, the question is equivalent to asking is \(x>3\) without any other information.

OpenStudy (anonymous):

woops. i was suppost to use Use the equation 8x 2 - 4xy + 5y 2 = 36 to find tan 2theta. (which is -4/3) could you help me answer it now?

OpenStudy (whpalmer4):

So \[\tan\,2\theta = -\frac{4}{3}\]and we are to determine which quadrant contains \(2\theta\)?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Is that going to be a positive value of \(\theta\) or a negative one?

OpenStudy (anonymous):

The options are quadrant 1, 2, or 4. I found theta (-26.56) so am i correct in saying its quadrant 4? i used theta= 1/2 tan^-1 (-4/3 )

OpenStudy (whpalmer4):

Well, I think quadrant IV is the right answer, but I'm not quite sure I see how you got \(\theta = -26.56\)

OpenStudy (anonymous):

I used tan 2 theta= B/ A-C i found that B/(A-C) is (-4/3) from the original equation. Then used theta = 1/2tan^-1 B/(A-C)

OpenStudy (whpalmer4):

Yes, I understand the formula you used, but I don't see how you get -26.56 as an answer. When I take the arctan of -4/3, I get -0.927295 radians.

OpenStudy (anonymous):

hmm.. i was using degrees and i probably messed up. (im using an iphone app) but i just replugged it in and got the same thing. so quadrant 4 is the right answer?

OpenStudy (anonymous):

Im probably going to ask more questions in this section